How many milliliters of 1.18 M NaOH must be added to 125 mL of 0.10 M NaH2PO4 to make a buffer solution with a pH of 6.90?
Answer #1

Consider a reaction of NaOH with NaH2PO4

NaH2PO4 + NaOH Na2HPO4

No of moles of NaH2PO4 = Molarity x volume of solution in L

= 0.1 mol / L x 0.125 L

= 0.0125 mol

Lets use ICE table

Concentration ( moles) NaH2PO4 NaOH Na2HPO4
Initial 0.0125 X 0
Change -X -X +X
Final 0.0125 - X 0.0 X

pH of buffer solution is calculated by using Henderson's equation as

pH = pKa + log [Na2HPO4 ] / [NaH2PO4]

Here pKa= 7.198

Therefore,

6.90 = 7.198 + log [Na2HPO4 ] / [NaH2PO4]

log [Na2HPO4 ] / [NaH2PO4] = 6.90 - 7.198 = - 0.298

[Na2HPO4 ] / [NaH2PO4] = 10 -0.298 = 0.504

[Na2HPO4 ] = 0.504 x [NaH2PO4]

X = 0.504 ( 0.0125 - X )

= 0.0063 - 0.504 X

X + 0.504 X = 0.0063

1.504 X = 0.0063

X = 0.0063 / 1.504 = 0.00419 = No of moles of NaOH

We have, Molarity = No of moles / Volume of solution in L

Volume of 1.18 M  solution in L = No of moles / Molarity

= 0.00419 mol / 1.18 mol / L

= 0.00355 L

= 3.55 ml

ANSWER: Add 3.55 ml of NaOH to the 125 ml of 0.1 M NaH2PO4